OpenStudy (anonymous):

i^i^i^i^i^e=0 i^i^i^i^i^i^e=1 i^i^i^i^i^i^i^e=i WHY

OpenStudy (perl):

i^i^i = -i

OpenStudy (kc_kennylau):

source please.

OpenStudy (perl):

actually im verifying this is true, one sec

OpenStudy (anonymous):

I meant ((i)^i)^i... not ((i^i)^i)^i... And I discovered this a while ago entering random queries in Wolfram Alpha, and then Google.

OpenStudy (kc_kennylau):

source please

OpenStudy (perl):

yes i just checked, five tetrations

OpenStudy (anonymous):

Well, I guess I can say the source is http://www.wolframalpha.com/input/?i=i%5Ei%5Ei%5Ei%5Ei%5Ei%5Ee partly.

OpenStudy (anonymous):

What's a tetration?

OpenStudy (perl):

2^2^2 is a tetration

OpenStudy (perl):

a tower of powers

OpenStudy (perl):

i = cos (pi/2 ) + i sin(pi/2) = e ^(i*pi/2) therefore i^i = [ e^(i*pi/2)]^i = e^ [ i*i*pi/2] = e^[ -pi/2]

OpenStudy (perl):

then you keep doing th same thing

OpenStudy (kainui):

\[i^i=(e^{i\frac{ \pi }{ 2 }} )^i =e^{-\frac{ \pi }{ 2 }}\]

OpenStudy (kainui):

Well @perl beat me I suppose.

OpenStudy (anonymous):

Where does that property come from, though?

OpenStudy (perl):

i^i^i = [[e^(i*pi/2)]^i)^i = (e^ [ i*i*pi/2])^i = (e^[ -pi/2])^i = e^[-Pi/2*i]

OpenStudy (perl):

which property?

OpenStudy (agent0smith):

\[\Large ((i)^i)^i\]etc right? You can use exponent rules to simplify some of it

OpenStudy (perl):